OpenStudy (anonymous):

DID I DO THIS CORRECTLY? Find the slope-intercept form for the line passing through (-3,3) and perpendicular to the line passing through (-7, 1/2) and (-4, 2/3). 2/3 - 1/2 / -4 - -7 = 1/6 / 3 = 1/18 slope of perpendicular line would be: 18 y-3=18(x--3) y-3=18(x+3) y-3=18x+54 The answer that I got: y=18x+57

OpenStudy (lalaly):

slope of perpindicular line is -18

OpenStudy (dumbcow):

If perpendicular slope is -18, then equation of line is y = -18x -51 y-3 = -18(x+3) y = -18x -51